OpenStudy (anonymous):

Find the exact value of each expression: a) tan^(-1) (1/sqrt3) b) sec^(-1) (2)

OpenStudy (anonymous):

\[\text{ArcTan}\left[\frac{1}{\sqrt{3}}\right]=\frac{\pi }{6}, \text{ ArcSec}[2]=\frac{\pi }{3} \]